Find the Derivative of the following function

  1. y = x2 cos x - 2xsinx - 2cosx
  2. x(cos2x + 3y) = ysinx
1
Expert's answer
2022-03-18T10:33:35-0400

"1.\\
y=x^2cosx-2xsinx-2cosx\\
y'=2xcosx+x^2(-sinx)-2sinx-\\-2xcosx-2(-sinx)=-x^2sinx\\
2.\\
x(cos2x+3y)=ysinx\\
y=\frac{xcos2x}{sinx-3x}\\
y'=\frac{(cos2x+x(-2sin2x))(sinx-3x)-xcos2x(cosx-3)}{(sinx-3x)^2}=\\
=\frac{(cos2x-2xsin2x)(sinx-3x)-xcos2x(cosx-3)}{(sinx-3x)^2}"